OpenStudy (pedro):

Solve 6 1/9 -3 1/3 . Show your work.

OpenStudy (anonymous):

when working with these kinds of problem, it would behoove you to change these mixed fractions into improper fractions

jhonyy9 (jhonyy9):

54+1 9+1 ----- - ------ = 55/9 - 10/3 = 9 3

jhonyy9 (jhonyy9):

= (55-30)/9 = 25/9